I am half my mothers age my father is 5 years older than my mother when you add up all our ages you get 100 years how old is my mother my father and myself
Answer by lwsshak3(11628)

I am half my mothers age my father is 5 years older than my mother when you add up all our ages you get 100 years how old is my mother my father and myself
***
let x=mother's age
x/2=my age
x+5=father's age
x+x/2+x+5=100
2.5x=95
x=38
x/2=19
x+5=43
mother's age=38
my age=19
father's age=43